1/ santa loses footing and slides down a frictionless, snowy roof that is tilted at an angle of 45 degress. If santa sildes 10 m before reaching the edge, what is his speed as he leaves the roof?
Answer #1

Santa's acceleration down the roof, a = 9.81*(cos45o) = 6.94 m/s2

Distance travelled down the roof, s = 10 m

Initial speed, u = 0

Speed as he leaves the roof, v = [u2 + 2as]1/2

=> v = [02 + (2 * 6.94 * 10)]1/2 = 11.78 m/s